OpenStudy (anabo):

FAN AND MEDAL Find the point that lies on the line described by the equation below. y - 4 = -2(x - 6)/ . (6, -4) B. (4, 6) C. (-12, 4) D. (6, 4)

OpenStudy (sunnnystrong):

Point Slope Form: \[(y-y1)=m(x-x1)\] ... what do you think? @Anabo

OpenStudy (triciaal):

what set of values ( from the options) will make the equation true?

OpenStudy (anabo):

in slope intercept form its y=-2x+16

OpenStudy (sunnnystrong):

@Anabo yep... so what do you think the answer is?

OpenStudy (anabo):

I'm still not completely sure but I'm thinking its A

OpenStudy (sunnnystrong):

hmmm.. If it was A-> Evaluate y @ x=6 y=-2(2)+16 y=12.

OpenStudy (anabo):

I'm really confused.

OpenStudy (sunnnystrong):

Oh shoot sorry XD My mistake Uhhh.. Not A.

OpenStudy (triciaal):

read the responses again then try some more

OpenStudy (sunnnystrong):

Evaluate y @ x=6 y=-2(6)+16 y=4.

OpenStudy (sunnnystrong):

What @triciaal said haha

OpenStudy (anabo):

So I have to pick whichever one is equivalent to y=-2x+16

OpenStudy (sunnnystrong):

You have to pick which ordered pair (x,y) satisfies the equation

OpenStudy (anabo):

is it D?

OpenStudy (sunnnystrong):

Yep (: